Jim can fill a pool carrying buckets of water in 30 minutes. Sue can do the same job in 45 minutes. Tony can do the same job in 1 1/2 hours.
How quickly can all three fill the pool together?

Answer by richwmiller(17219)

In one minute, Jim can do 1/30=3/90
In one minute, Sue can do 1/45=2/90
In one minute,Tony can do 1/90=1/90
3/90t+2/90t+1/90t=1
(6/90)*t=1
Solve for t
We would expect the job to be done by all three in much less than 30 minutes since Jim can do it alone in 30 minutes